OpenStudy (anonymous):

Five weeks after the introduction of a new toy, the manufacturer found that the demand for the toy was modeled by f(x)=x^2+5x+4/x^2+6x-7 - 1/2 , while the company’s supply capability was modeled by g(x)= -0.89x^2+850x-50/800x+70.5 . Here x is the number of weeks the toy has been on the market, and the values of f(x) and g(x) are in millions of toy units. About how many weeks after the toy was introduced did supply equal demand? Round to the nearest whole number of weeks.

OpenStudy (anonymous):

hi @madiluv9429 I'm reading your question and will try to help out with it

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

let me do some work real fast… and from there see if i can explain it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[f(x)=\frac{ x^{2}+5x+4 }{ x^{2}+6x-7 }-\frac{ 1 }{ 2 }\] is that what your f(x) should look like

OpenStudy (anonymous):

the one half is actually 1/x sorry

OpenStudy (anonymous):

tough question, but when i graphed both equations i see the number of weeks would equal 3.37

OpenStudy (anonymous):

or the nearest whole =3weeks

OpenStudy (anonymous):

i'm not sure how to explain, but i can post a pic of my graphing

OpenStudy (anonymous):

its very confusing because i got the same answer but my answers are A. 74 weeks B. 80 weeks C. 76 weeks

OpenStudy (anonymous):

ok i see now

OpenStudy (anonymous):

i zoomed way out on my graph and saw another intersection

OpenStudy (anonymous):

@ 75.7 weeks or 76

OpenStudy (anonymous):

ohkay can you help me on the other questions i got wrong

OpenStudy (anonymous):

iam willing to give it a shot

OpenStudy (anonymous):

okay give me two secs

OpenStudy (anonymous):

np

OpenStudy (anonymous):

his graph of a function is a translation of y=2/x . What is an equation for the function?

OpenStudy (anonymous):

sec

OpenStudy (anonymous):

the answers are A.y=2/x-3 -5 B. y=2/x-5 -3 C. y=2/x-5 +3

OpenStudy (anonymous):

y=(2/x-5))-3 or b

OpenStudy (anonymous):

do you know how to do transformations

OpenStudy (anonymous):

I'm sure you do

OpenStudy (anonymous):

yes i do but this is suppose to be algebra 2 and non of this is in my books so i am pretty confused and this is my last class til i graduate so i need to pass

OpenStudy (anonymous):

Simplify the difference. \[n ^{2}-10n+24/n ^{2}-13n+42 - 9/n-7\]

OpenStudy (anonymous):

\[\frac{ n^{2}-10n+24 }{ n^{2}-13n+42 }-\frac{ 9 }{ n-7 }\] is this how it looks

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k now i will factor the numerator and denominator of the left side for….sec

OpenStudy (anonymous):

\[\frac{ (n-4)(n-6) }{ (n-6)(n-7) }-\frac{ 9 }{ (n-7) }\]

OpenStudy (anonymous):

ok so you took away the exponents

OpenStudy (anonymous):

yes now we cancel like terms of the num and denim

OpenStudy (anonymous):

denom*

OpenStudy (anonymous):

so we cancel out n-4 and n-6 and n-6 and n-7

OpenStudy (anonymous):

just looking at the left side, we will cancel the (n-6) from top and bottom

OpenStudy (anonymous):

ok because theres n-6 onn top and bottom?

OpenStudy (anonymous):

yesm and that will leave us with \[\frac{ (n-4) }{ (n-7) }-\frac{ 9 }{ (n-7) }\]

OpenStudy (anonymous):

ok now we cancel out the n-7 aswell?

OpenStudy (anonymous):

nope that will remain, that is our common denominator

OpenStudy (anonymous):

\[\frac{ (n-4)-9 }{ (n-7) }\]

OpenStudy (anonymous):

so we just combined the n-7 and kept the 9 as a negative

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can still write it as before \[\frac{ (n-4) }{ (n-7) }-\frac{ 9 }{ (n-7) }\]

OpenStudy (anonymous):

because both (n-4) and 9 are being divided by the same denominator (n-7)

OpenStudy (anonymous):

they just wanted it simplified! thats all

OpenStudy (anonymous):

im not able to see them

OpenStudy (anonymous):

oh wait weird

OpenStudy (anonymous):

those are the given answers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i want to say a because it has the same denominator i dunno where the 13 came from one sec

OpenStudy (anonymous):

was there a b answer

OpenStudy (anonymous):

you posted the three answers above a, c, and d

OpenStudy (anonymous):

i put be and i got it wrong

OpenStudy (anonymous):

what was answer b?

OpenStudy (anonymous):

you didn't post it earlier

OpenStudy (anonymous):

n-4/n-7

OpenStudy (anonymous):

what about the 9? (n-4)-9/(n-7)

OpenStudy (anonymous):

there was no 9

OpenStudy (anonymous):

did they give you the right answer?

OpenStudy (anonymous):

no i have to figure out on my own

OpenStudy (anonymous):

they allow multiple chances to answer?

OpenStudy (anonymous):

i say its a because it shows the same denominator, but idk where the 13 came from

OpenStudy (anonymous):

yes i do online and so everything except the final is multiple choice answers

OpenStudy (anonymous):

9+4 is 13 but i don't see how that would be legit

OpenStudy (anonymous):

i dont know thats why im so confused

OpenStudy (anonymous):

or -4- -9= -13

OpenStudy (anonymous):

oh i know whats up

OpenStudy (anonymous):

should have just factored the left side denim first to find the common denim

OpenStudy (anonymous):

@Mertsj might be able to help

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Simplify the difference. n^2−10n+24/n^2−13n+42−9/n−7 is the initial question @merts

OpenStudy (mertsj):

The problem you have to solve is f(x)=g(x)

OpenStudy (anonymous):

@Mertsj she has a new question

OpenStudy (anonymous):

simplifying \[\frac{ n^{2}-10n+24 }{ n^{2}-13n+42 }-\frac{ 9 }{ n-7 }\]

OpenStudy (anonymous):

then factored, had a common denominator, n got stuck

OpenStudy (mertsj):

\[\frac{n-4}{n-7}-\frac{9}{n-7}=\frac{n-4-9}{n-7}=\frac{n-13}{n-7}\]

OpenStudy (anonymous):

i was right but i didn't know where the -13 came from

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

@madiluv9429 it was A :)

OpenStudy (anonymous):

okay thank you i need help on other questions as well

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Simplify the sum 4/m+9 + 5/m^2-81

OpenStudy (anonymous):

\[\frac{ 4 }{ m+9 }+\frac{ 5 }{ m ^{2}-81 }\] like this?

OpenStudy (anonymous):

@madiluv9429

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok \[\frac{ 4 }{ m+9 }+\frac{ 5 }{ (m+9)(m-9) }\]

OpenStudy (anonymous):

you find a common number that will go into 81 and 9 and that is 9 is that right

OpenStudy (anonymous):

yep when m^2-81 factored is (m+9)(m-9)

OpenStudy (anonymous):

ohhh ohkay than what

OpenStudy (anonymous):

i think it is just 9/m+9